Write as a mixed number.140 1/2%
Answer by stanbon(57290)

Write as a mixed number.140 1/2%
------------
140 1/2%
= 140.5/100
=1.405
=1+(405/1000)
=1+81/200
= 1 81/200
Cheers,
Stan H.